### 16 Apr 2026 v3

# Generalized saddle-node ghosts and their composite structures in dynamical systems

Abstract
The study of dynamical systems has long focused on the characterization of their asymptotic dynamics such as fixed points, limit cycles and other types of attractors and how these invariant sets change their properties as systems parameters change. More recently, however, the importance of transient dynamics, especially of long transients and sequential transitions between them, has been increasingly recognized in various fields including ecology, neuroscience, and cell biology. Among several possible origins of long transients, ghost attractors have received particular attention due to interesting dynamical properties in non-autonomous settings, new theoretical developments, and an increasing number of systems that empirically exhibit dynamics consistent with ghost attractors. Despite this growing interest in transient dynamics generally and ghost attractors in particular, there are significantly fewer theoretical concepts and software tools available to researchers to classify and characterize their underlying mechanisms compared to asymptotic dynamics. To address this gap, we generalize saddle-nodes to account for higher-dimensional center manifolds and provide a definition for their ghost attractors. We then introduce algorithms to specifically identify and characterize ghost attractors and their composite structures such as ghost channels and ghost cycles and show how these concepts and algorithms can be used to gain new insights into the transient dynamics of a wide range of system models focusing on living systems, allowing, e.g., to describe bifurcations of ghosts. The algorithms are implemented in Python and available as PyGhostID, a user-friendly open-source software package.
Keywords Ghost attractors $\cdot$ Bottlenecks $\cdot$ Ghost channels $\cdot$ Ghost cycles $\cdot$ Metastability $\cdot$ Transient dynamics $\cdot$ Saddle-node bifurcation $\cdot$ SNIC bifurcation $\cdot$ Dynamical Systems $\cdot$ Python package
references_clean.bib]
1 Introduction
Much of dynamical systems research focuses on understanding a system of interest through the lens of asymptotic dynamics, i.e., by studying how invariant sets such as fixed points, limit cycles and other types of attractors change as systems parameters are varied. Examples of this approach include tipping of climate- and ecosystems into undesirable states via passing through saddle-node bifurcations [59, 32, 60, 73, 41], the birth of the heart beat as emergence of a limit cycle [34] and its degeneration into cardiac arrhythmia as transition to chaos [24, 72, 58], or cellular decisions as the selection of distinct attractors of multi-stable molecular regulatory networks [76, 49, 11]. Although studying attractors and their transitions between them has indisputably generated crucial insights into innumerable systems across the sciences, there are limitations to this approach. Indeed, many real-world systems often do not exhibit true asymptotic dynamics, e.g., due to frequent influences from their surroundings that render them non-autonomous. Some forms of cardiac arrhythmia, for instance, may best be described by transients that occur during transitions between steady states triggered by a sudden input signal rather than as asymptotic dynamics [75]. In other cases, systems seemingly reside in a stable state before suddenly shifting towards a qualitatively different regime even in absence of any change of parameters or external inputs to the system, challenging the conceptualization of such dynamics as attractors. The importance of transient dynamics, especially of long transients - or metastable states - and sequential transitions between them, is thus increasingly recognized in various fields, including ecology [27, 48], neuroscience [54, 25] and cell biology [20, 35, 38].
Several mechanisms have been put forth to explain the emergence of long transients (see, e.g., [27] and [54] for reviews). Historically the first and mathematically best understood one among these is the explicit separation of timescales, resulting in what are called today slow-fast or multi-timescale systems [66, 70, 39, 7]. In such systems, one or more processes take place on a timescale that typically differs by an order of magnitude or more and is often accounted for with a small parameter $\epsilon>0$, scaling the rate of change for one of the system variables. Using geometric singular perturbation theory, the slow dynamics in such systems can be studied as the flow on the so-called critical manifold in the limit of $\epsilon=0$, the results of which can be shown to hold for the slow manifold at $\epsilon>0$ due to Fenichel theory [21, 22]. While not all slow-fast or multi-timescale systems give rise to long transients, the cyclical visitation of slow manifolds in many oscillatory systems such as the FitzHugh-Nagumo model or the van der Pol oscillator naturally gives rise to alternating sequences of long transients. A different mechanism frequently invoked in ecology and neuroscience is the visitation of saddles by the system’s phase space trajectory, either in the form of saddle "crawl-bys" [27] or heteroclinic channels/cycles [43], causing the system to slow down in the vicinity of the visited saddle fixed point. Heteroclinic channels, cycles and networks are particularly suited for modeling sequential transitions between multiple transient states and have found applications, e.g., in neuroscience [52, 1] and robotic control [30].
Another notable mechanism, and the focus of this paper, is the slowing down due to a bottleneck or ghost [64, 65] when a system is close to a saddle-node (SN) bifurcation. From a geometrical perspective, the slow dynamics results from a plateau in the potential landscape of the system which is a remnant of a former attractor close-by in parameter space [63, 38]. Ghosts are a well-known theoretical mechanism for long transients underpinned by a growing body of empirical evidence in many systems including shallow-water lakes [68], neuronal dynamics [33], and cells [62, 53, 45]. Being close to bifurcation, ghosts are particularly interesting in the context of non-autonomous systems, i.e., when system parameters or inputs are explicitly time-varying, often resulting in quite different responses compared to pure slow-fast systems [36]. This feature has also been shown to improve information processing capabilities in biological systems [63, 47, 13]. More recently, we showed that multiple ghosts can be connected to each other to form larger composite structures, which we termed ghost channels and ghost cycles [38], providing an alternative description of sequential long transients. The trapping times in the presence of noise and the response of these objects to periodic forcing are quite different from heteroclinic structures [38, 51] and slow-fast systems [36], respectively.
Despite this variety of mechanisms, the theoretical and the computational tools available to researchers for characterizing transient dynamics are sparse compared to the tools available for asymptotic dynamics. While the lack of theoretical tools has sparked several promising new research avenues [40, 28, 77], the lack of computational tools persists. While many powerful and widely used software solutions exist, e.g., for the characterization of attractors and their bifurcations [17, 18, 16, 69, 14], the few algorithms and software solutions available to characterize transient dynamics (e.g., for numerical approximation of slow-manifolds via Computational Singular Perturbation [42]) are often rather domain specific or buried in sparsely documented research code. Thus, almost no widely applicable, user-friendly plug-and-play solutions for the characterization of transient dynamics in arbitrary dynamical systems exist to date, requiring researchers to resort to time-consuming manual analysis.
In this article, we address this gap for one of the possible mechanisms underlying long transients and sequential transitions between them: ghost states and their composite structures. By giving a novel generalized definition of ghost states, we provide a new theoretical framework that allows to better describe ghosts as phase space objects and phenomena associated with ghosts, many of which have not been reported before. Based on this framework, we introduce a series of novel algorithms for identifying and characterizing ghost states and their composite structures (e.g., ghost channels [38]) along system trajectories, enabling us to obtain new insights into the transient dynamics, e.g., of neuronal and gene regulatory network models and to discover novel transient dynamics phenomena such as bifurcations between ghosts.
The remainder of this article is organized as follows. In subsection 2.1, we review the basic concepts required to understand ghost states and their properties before giving a new formal definition that captures these features and generalize it to higher dimensional ghosts. Following the outline of the basic framework, in subsection 2.2 we describe the main algorithm of this paper used to identify and characterize ghost states from trajectories and how it is embedded in the PyGhostID package. In subsection 2.3, we validate the algorithm, showing that GhostID specifically identifies ghost states but not saddles or slow-fast dynamics as origins of long transients. In subsection 2.4, we show how our concepts and algorithms can be used to study various dynamical systems models with a focus on living systems. We conclude in section 3 by discussing current limitations of our approach, implications of the findings presented here, and avenues for future research.
2 Results
2.1 Theoretical framework
Before considering the non-asymptotic dynamics of ghost states, we briefly review some basic concepts about how the flow of a dynamical system is organized by fixed points (equilibria) as some of the most basic objects that govern a system’s asymptotic dynamics. From a geometric perspective, fixed points are important because they direct the flow along defined directions and thereby determine a trajectory’s fate. Given an autonomous dynamical system $\dot{x}=f(x,\rho)$ with parameters $\rho$, a fixed point is simply defined as any $x^{*}$ with $f(x^{*},\rho)=0$. Different types of fixed points are classified by how the system’s dynamics is organized in their vicinity. Stable fixed points (also called attractors or sinks) attract all trajectories from their neighborhood, whereas unstable fixed points (also called repellers or sources) direct trajectories to move away from the fixed point. As a simple example, consider the following system for $\mu<0$:
$\dot{x}=f(x,\mu)=\mu+x^{2}.$ (1)
Visualizing the dynamics by plotting $\dot{x}$ vs $x$, we immediately see that the system has two fixed points, one at $x_{s}^{*}=-\sqrt{\mu}$ and one at $x_{u}^{*}=\sqrt{\mu}$ (Figure 1a). Inferring the direction of flow by considering the value of $\dot{x}$ in the vicinity of $x_{s}^{*}$ (black arrowheads in Figure 1a), we can see that trajectories will move towards $x_{s}^{*}$, hence $x_{s}^{*}$ is a stable fixed point. By the same kind of reasoning, we find that $x_{u}^{*}$ is unstable. In higher dimensions, sinks and repellers exist, too, but are complemented by saddles, defined to have one repelling and one attracting direction on the plane, or $k$, $0<k<n$, repelling and nkn-k attracting directions in $n$ dimensions, respectively (Figure 1b).
The notion of stability and the classification of fixed points can be made precise by linear stability analysis, in which a system is linearized around the fixed point and classified by the eigenvalues $\lambda_{1},\dots,\lambda_{n}$ of the Jacobian matrix $D_{x}f(x^{*})$ at the fixed point: a fixed point $x^{*}$ is a repeller if $\forall i:\textnormal{re}(\lambda_{i})>0$, a sink if $\forall i:\textnormal{re}(\lambda_{i})<0$ or a saddle if there are $k$ eigenvalues with $\textnormal{re}(\lambda_{i})>0$ and nkn-k eigenvalues with $\textnormal{re}(\lambda_{i})<0$, respectively. Fixed points with $\forall i:\textnormal{re}(\lambda_{i})\neq 0$ are called hyperbolic. In this study, however, we will be primarily concerned with fixed points that have at least one zero eigenvalue, i.e., with non-hyperbolic fixed points. Non-hyperbolic fixed points commonly arise at bifurcations and are quite different from their hyperbolic counterparts: their stability cannot easily be determined via linear stability analysis and they are typically quite fragile (structurally unstable) as the topology of the phase portrait is altered by arbitrarily small perturbations to the vector field. While non-hyperbolic fixed points come in different varieties, we will focus next on saddle-node fixed points (not to be confused with hyperbolic saddles) and their ghosts as organizing centers of a system’s non-asymptotic dynamics.
To gain intuition about saddle-nodes, we return to Equation 1. As we increase $\mu$, $x_{s}^{*}$ and $x_{u}^{*}$ move closer and closer together until they merge into a single fixed-point at $\mu=0$ (Figure 1c), which is called a one-dimensional saddle-node - the namesake of the SN-bifurcation. The eigenvalue of the Jacobian of a one-dimensional first-order ODE is equivalent to the first derivative evaluated at the fixed point, i.e., in this case $\lambda=f^{\prime}(0,0)=0$, confirming that our saddle-node is non-hyperbolic. Since $f(x,0)>0$ for both $x<0$ and $x>0$, all points on the negative half-line will approach $x^{*}=0$ as $t\to\infty$, whereas all points on the positive half-line will move away from it, such saddle-nodes are sometimes called semi-stable. In this sense, the saddle-node inherited the negative half-line as a basin of attraction from the stable fixed point, and a repelling positive half-line from the unstable fixed point. Moving on to higher dimensions, we can see that various types of saddle-nodes can be constructed by pairing one or more semi-stable directions with stable and unstable directions that correspond to a system’s center, stable and unstable eigendirections, respectively (Figure 1d). This picture makes it furthermore clear that in higher dimensions, saddle-nodes exist whose eigenspace is identical with their center subspace and that many saddle-nodes have unstable directions, repelling all trajectories that do not begin exactly on a stable or center manifold. To describe these saddle-nodes formally, we begin by generalizing the definition of a saddle-node equilibrium of type $k$ [3, 4] to account for semi-simple eigenvalues zero:
Figure 1: Organization of phase space flow by fixed points and saddle-node ghosts. a, b, hyperbolic fixed points in dimensions 1-3. a, sink and repeller for flows on the line (Equation 1) for which stability of fixed points can be visually assessed by considering $\dot{x}$ versus $x$. c,d, non-hyperbolic saddle-nodes in dimensions 1-3. c, saddle-node at the bifurcation point for flows on the line. Since $\dot{x}>0$ for all $x\neq 0$, the saddle-node at $x=0$ is semi-stable: the negative half-line is attracted to the fixed point, the positive half-line moves away from it. d, saddle-nodes of type $j,k$ in dimensions 2 and 3. Black arrows indicate eigenvectors of the stable/unstable subspaces, magenta arrows indicate eigenvectors of the center subspace. e,f, ghosts for various types of saddle-nodes in dimensions 1-3. e, ghost for flows on the line. While $\dot{x}>0$ for all $x$, the general direction of the flow is identical as for the saddle-node and low values for $\dot{x}$ at $x\approx 0$ (bottleneck) lead to a long transient (inset) which together constitutes finite-time attraction. e, ghosts of type $j,k$ saddle-nodes in dimensions 2 and 3. Dotted arrows indicate flow directions at the ghost. Note how despite the absence of fixed points the phase portrait maintains high similarity to the parental saddle-node.
Definition 2.1 (Saddle-node equilibrium and bifurcation point of type $j,k$).
Let $\dot{x}=f(x,\rho_{\textnormal{crit}})$, $x\in\mathbb{R}^{n}$, $f\in\mathcal{C}^{\omega}$, be an autonomous dynamical system with parameters $\rho_{\textnormal{crit}}=(\rho_{1},\dots,\rho_{m})^{T}\in\mathbb{R}^{m}$. We say $x^{*}$ fulfilling $f(x^{*})=0$ is a saddle-node equilibrium of type $j,k$ and $(x^{*},\rho_{\textnormal{crit}})$ a saddle-node bifurcation point of type $j,k$ if the following holds true:
  1. (SN1)
    $D_{x}f(x^{*},\rho_{\textnormal{crit}})$ has a semi-simple eigenvalue 0 of algebraic and geometric multiplicity $j$, with corresponding linearly independent eigenvectors $v_{1},\dots,v_{j}$ and $w_{1},\dots,w_{j}$ to the right and left, respectively, that vary continuously with $(\rho_{1},\dots,\rho_{m})^{T}$.
  2. (SN2)
    For each $i$ with $1\leq i\leq j$ there exists $l$ with $1\leq l\leq m$ with $w_{i}((\partial f/\partial\rho_{l})(x^{*},\rho_{\textnormal{crit}}))\neq 0$.
  3. (SN3)
    for every non-zero $u\in\ker(D_{x}f(x^{*},\rho_{\textnormal{crit}}))$, there exists $i$ with $1\leq i\leq j$ such that $w_{i}(D_{x}^{2}f(x^{*},\rho_{\textnormal{crit}})(u,u))\neq 0$.
  4. (SN4)
    $D_{x}f(x^{*},\rho_{\textnormal{crit}})$ has $k$ eigenvalues $\lambda^{+}_{1},\dots,\lambda^{+}_{k}$ with positive real parts, and njkn-j-k eigenvalues $\lambda^{-}_{1},\dots,\lambda^{-}_{n-j-k}$ with negative real parts.
For $j=1$, Definition 2.1 reduces to the definitions given in [3, 4], but for $j>1$ captures the other saddle-nodes shown in Figure 1d: (SN1) ensures a center subspace space spanned by $j$ independent eigendirections, (SN2) means that for each center direction, there is a parameter that upon perturbation leads to an unfolding of the saddle-node, i.e. saddle-nodes of type $j,k$ are structurally unstable. Note that since this can include different parameters for different center directions, the saddle-node of type $j,k$ has a codimension of $\leq j$. (SN3) guarantees quadratic non-degeneracy for each non-zero center direction. (SN4) simply describes the dimension of the saddle-nodes’ stable and unstable subspaces.
Having defined saddle-nodes of type $j,k$, we next will consider how their ghost influences the non-asymptotic dynamics of a system after a SN-bifurcation occurs. To do so, we return again to Equation 1, this time for a small $\mu>0$ (Figure 1e). Since $\dot{x}=\mu+x^{2}>0$ for $\mu>0$, Equation 1 has no fixed points anymore, i.e., no invariant objects can govern the system’s dynamics. However, while $\dot{x}>0$, $\mu$ is small and so for small $|x|$, $\dot{x}$ becomes small, too, and any trajectory starting on the negative half-line spends a long time in the vicinity of 0 before eventually escaping towards infinity (cf. inset in Figure 1e). Notice how the phase portrait of the ghost, except for the vanished fixed point, does not differ much from that of its saddle-node at $\mu$, in particular the direction of the flow remains the same. Indeed, this simple example already shows many features that are associated with ghosts: slow dynamics, non-invariance and a finite-time sense of attraction. However, as saddle-nodes can become more complex in higher dimensions, so do their ghosts. Consider the following system:
$\dot{x}_{i}=\begin{cases}\mu_{i}+x_{i}^{2},&1\leq i\leq j,\\[6.0pt] x_{i},&j+1\leq i\leq j+k,\\[6.0pt] -x_{i},&j+k+1\leq i\leq n\end{cases}$ (2)
where integers $n,j,k$ describe the system’s dimension, the number of dimensions that follow normal form dynamics, and the number of attracting and repelling directions, respectively, and $\mu_{i},1\leq i\leq j$ are the parameters. By design, subsection 1.7 describes saddle-nodes of type $j,k$ in $\mathbb{R}^{n}$ for any $n$ at $\mu_{1}=\dots=\mu_{j}=0$. For $\mu_{1}=\dots=\mu_{j}>0$, we see ghosts for each saddle-node of type $j,k$ emerging, each of whose phase portrait closely resembles the corresponding saddle-node (Figure 1f). Ghosts of saddle-nodes of type $1,0$ are the ‘standard’ ghosts commonly found in the literature: funneling trajectories from a larger area of the phase space, slowing them down and ejecting them into a single direction. In contrast, ghosts of type $j,k$ saddle-nodes with $k>0$ repell most trajectories, defying the notion of attraction. Ghosts of type $j,0$ saddle-nodes, $j>1$, inherit the phase-portrait organized by the crossing center eigendirections of the saddle-nodes, collecting and bundling trajectories from one region of the phase space, slowing them down locally, but then ejecting them into more than one directions (cf. s.-n. ghost type $2,0$ in Figure 1f), in that sense making ghosts of $j,0$ saddle-nodes, $j>1$, genuinely higher-dimensional. While this phenomenological description shows that our standard notions do not apply to all ghosts, we still have not defined formally what a ghost is. Perhaps owing due to their non-invariant and transient nature, there is, in fact, no single and generally accepted formal definition of ghosts in dynamical systems yet. The recent definition by Zheng et al. [77], for example, considers ghosts as the non-zero minima of user-defined cost functions to capture the slow dynamics and enable their continuation across parameters, whereas our previous definition [38] accounts for their non-invariance and the existence of a basin of attraction to define composite structures made up from multiple ghosts in phase space. However, neither definition captures all of the above described features and is sufficient for the purpose of this study. In particular, local minima of some of the cost functions described in [77], as we will see later, also occur in systems with long transients that are not due to ghosts and neither [77] nor [38] account for the dimensionality of a ghost. We thus propose the following new definition and terminology:
Definition 2.2 (Ghost of type $j,k$ saddle-node).
Let $\dot{x}=f(x,\rho)$, $x\in\mathbb{R}^{n}$, $f\in\mathcal{C}^{\omega}$, be an autonomous dynamical system with parameters $\rho=(\rho_{1},\dots,\rho_{m})^{T}\in\mathbb{R}^{m}$ and $\mathcal{A}\in\mathbb{R}^{n}$ a closed bounded set. We say the system has a ghost of a type $j,k$ saddle-node at $x_{g}\in\mathcal{A}$ if the following holds:
  1. (G1)
[…]
  2. (G2)
    $\mathcal{A}$ contains no semi-trajectories in forward time (non-invariance)
  3. (G3)
[…]
We call $j$ the dimension of the ghost and say ghost $x_{g}$ is attracting if its limiting saddle-node is of type $j,0$ and non-attracting otherwise.
While we have not explicitly stated how the set $\mathcal{A}$ is to be defined, we find in practice a small neighborhood of a potential ghost to be sufficient. (G1) formally captures the slow dynamics that result from the flat quasi-potential landscape due to the coalescence of equilibria at the saddle-node [62, 38, 37]. The auxiliary function $Q$ is equivalent to the cost-function $J$ used in [77]. (G2) simply states the non-invariance of ghosts, i.e., any trajectory visiting the neighborhood of a ghost must eventually leave this neighborhood again. While (G1-G2) are necessary conditions for ghosts, they are likely met by most long transients. (G3) is the core of the definition and states that each ghost has a closest saddle-node of type $j,k$ in parameter space from which it inherits the organization of the phase portrait and which defines the ghost’s dimension. It also excludes that a ghost can be a ghost of multiple saddle-nodes and will be important in the context of different ways of unfolding a type $j,k$ saddle-node. Although the notion of attraction is only implicit (as it is not necessary for the purpose of this study), definition 2.2 accounts for all features observed for the ghosts highlighted above.
2.2 The GhostID algorithm
Having formally defined what we mean by a ghost brings us a step closer to an algorithm that can identify and characterize ghosts. While (G1) and (G2) are straightforward to test, searching the parameter space for parental saddle-nodes is not a feasible strategy. We thus need move on from (G3) to an algorithmically identifiable criterion. For this, we will rely on our previous observation [38] that all ghosts we examined so far exhibit a linear spatial distribution of instantaneous eigenvalues ranging from negative to positive along the former center direction of the saddle-node. Indeed, we find the same to be true for higher dimensional ghosts (e.g., for system 1.7 for $n=j=2,k=0$, cf. Supplementary Figure 1). This leads us to formulate the following conjecture:
Conjecture 2.1 (Eigenvalue spectrum of type $j,k$ saddle-node ghosts).
Let $\dot{x}=f(x,\rho)$, $x\in\mathbb{R}^{n}$, $f\in\mathcal{C}^{\omega}$, be an autonomous dynamical system with parameters $\rho=(\rho_{1},\dots,\rho_{m})^{T}\in\mathbb{R}^{m}$ for which $x_{g}\in\mathcal{A}$ is a ghost of a type $j,k$ saddle-node $x_{sn}$ at $\rho_{\textnormal{crit}}$. For each $i$ with $1\leq i\leq j$, there exists a unit vector $u_{i}\in\mathbb{R}^{n}$ such that in the neighborhood of $x_{g}$, the eigenvalues $\lambda_{i}$ of $D_{x}f(x,\rho)$ along the line $\{x\in\mathcal{A}\ |\ x_{g}+tu_{i},\ t\in\mathbb{R}\}$ vary continuously and change sign from negative to positive with $\lambda_{i}(t)\approx ct$, $c\in\mathbb{R}$, to leading order.
Conjecture 2.1 states that the instantaneous eigenvalue distributions along the former center directions of the saddle-node still persist at the ghost, which is what we will use as algorithmic criterion to evaluate if a trajectory passes nearby a ghost. Although we were not able to provide a proof here, we strongly suspect conjecture 2.1 follows from definition 2.1 and 2.2 and find empirically that such a distribution exists in all systems exhibiting a ghost state resulting from a SN-bifurcation. While such a spatial distribution of instantaneous eigenvalues might potentially result from other mechanisms, too, and hence does not strictly guarantee the presence of a ghost, we empirically find it to be a very strong indicator based on numerous tested systems with different mechanisms underlying long transients.
Figure 2: Illustration of the main steps of how ghosts are identified via the GhostID algorithm.
Following the outlined framework, we propose a simple algorithm, which we term GhostID, that identifies ghosts along a trajectory by evaluating criteria (G1), (G2) and the changes in instantaneous eigenvalues from conjecture 2.1. The basic steps of this algorithm, illustrated in Figure 2, are as follows: The first step of the algorithm is to find the locally slowest points along a trajectory, corresponding to evaluating a trajectory proxy for criterion (G1). In the second step, the algorithm evaluates if the trajectory leaves the $\varepsilon$-neighborhood of a locally slowest point, testing for non-invariance (G2). The third and final step consists of testing whether one or more of the instantaneous eigenvalues change from negative to positive along the trajectory segment within the $\varepsilon$-neighborhood from the second step. Thus, starting from a slow point, if both non-invariance and the eigenvalue criterion are satisfied, we consider the long transient to be caused by a ghost, with the number of eigenvalues changing sign being considered by us as the ghost’s dimension. A full technical description of the algorithm, its hyper-parameters and some rationales for choosing their values them can be found in Supplementary section 1.1 and Supplementary Figure 2a-b.
Besides the GhostID algorithm, we developed three further algorithms that are build on GhostID and represent core functionalities of the PyGhostID package (Supplementary Figure 2c): ghostID_\_phaseSpaceSample, an integrated simulation and evaluation version of the algorithm that uses Latin hypercube sampling and parallel processing to sample many trajectories within a phase space region for ghosts, ghost_\_connections, which allows to reconstruct ghost channels, ghost cycles [38] and more complex composite structures of ghosts from one or more sequences of ghosts identified by GhostID, and track_\_ghost_\_branch, which, inspired by [77], allows to track an identified ghost state as system parameters are varied. While a more detailed description of these functions can be found in Supplementary section 1.2 and the PyGhostID documentation, we provide examples of their application in subsection 2.4.
2.3 Numerical validation of the GhostID algorithm
Before applying GhostID to obtain new insights, we first perform a series of numerical validations to ensure the algorithm correctly identifies ghosts in models from various disciplines and does not identify long transients as ghost that result from other mechanisms such as slow manifolds from slow-fast systems or saddle crawl-bys.
To begin our validation, let us consider the following model of ecological interactions between corals and macroalgae by Bieg et al. [8]:
$\dot{C}$ $=C\left(r(1-CN_{t})B-m-aM\left(\frac{N_{t}}{N_{0}+N_{t}}\right)\right)$ (3)
$\dot{M}$ $=M\left(aC\left(\frac{N_{t}}{N_{0}+N_{t}}\right)-\frac{g}{M+B}+\gamma B\right)$
$B$ $=1-C-M,$
where $C$, $M$ and $B$ represent corals, macroalgae and benthic r-strategists (groups of algae that colonize disturbed reefs on a faster timescale than macroalgae), and parameters $a,b,r$ and $\gamma$ represent (over-)growth rates, $g$ the grazing rate by herbivores, and $m$ is the mortality rate of corals. This model has been proposed to be organized close to a saddle-node bifurcation where it exhibits a ghost attractor to explain how coral reef recovery may be hampered by repeated climate or human driven perturbations that prevent trajectories escaping a long transient of the algae-dominated state [8]. Applying GhostID to a trajectory from this regime correctly identifies a ghost between the system’s nullclines (Figure 3a, left). Considering the $pQ$ time series reveals that the algorithm interestingly identified two $Q$-minima along the trajectory (Figure 3a, middle). Once the trajectory reaches the system’s stable fixed point, $pQ(t)$ appears to fluctuate irregularly, leading to additional $Q$-minima. However, these irregular movements are mere numerical noise which can be reduced by lowering the numerical tolerances of the integration procedure (Supplementary Figure 4) and have been ignored by GhostID by requiring identified peaks to have a minimum width of 5050dt. The eigenvalues along the trajectory segment around the first identified $Q$-minimum are consistent with the system visiting a slow region around the saddle on the y-axis and are thus not identified as a ghost, whereas on the segment of the second minimum eigenvalue $\lambda_{2}$ changes sign from negative to positive, leading to the algorithm identifying a ghost (Figure 3a, right). Additional examples of GhostID correctly identifying known ghosts in models from cellular biochemistry, gene regulatory networks, and condensed matter physics can be found in Supplementary section 1.4.1 and Supplementary Figure 5. Lastly, while we are not aware of higher-dimensional or non-attracting ghosts being described in the literature, GhostID correctly identifies these ghosts and their corresponding dimensions from our example system 1.7 (Supplementary Figure 6).
Having confirmed that GhostID correctly identifies previously characterized ghost states in models of various real-world phenomena, we next test if GhostID correctly ignores long transients resulting from other phenomena. First, we test GhostID on trajectories with long transients resulting from passing near saddles. The first system we consider for this task is a simple ecological model from Hastings et al. [27]:
$\dot{N}=\alpha N\left(1-\frac{N}{K}\right)-\gamma\frac{NP}{N+h},\quad\dot{P}=\epsilon\left(v\gamma\frac{NP}{N+h}-mP\right),$ (4)
where $N$ denotes the prey species, $P$ the predators, and $\alpha$ is the growth rate of the prey population, $K$ its carrying capacity, $\gamma$ the predation rate, $h$ the half-saturation constant for predation, $v$ is a scaling factor of how predation influences predator reproduction, $m$ is the mortality rate of predators and $\epsilon$ the timescale separation between predator and prey dynamics. For the parameter regime depicted in Figure 3b (left), the system exhibits oscillatory dynamics in which the orbit is forced to pass close to two different saddles (saddle crawl-by), causing the trajectory to slow down in their vicinity. However, the eigenvalues along the trajectory segments of the corresponding $Q$-minima do not cross 0, hence the saddle crawl-bys are not identified as ghosts by our algorithm (Figure 3b, middle and right). Another example involving a heteroclinic cycle can be found in Supplementary section 1.4.2 and Supplementary Figure 7. Since saddles are hyperbolic fixed points by definition, it is easy to prove that within a sufficiently small neighborhood of the fixed point, the eigenvalues along a trajectory in this neighborhood can never change sign (Supplementary section 1.5) and thus will not be identified as ghosts by our algorithm.
The next important mechanism to test our algorithm against are long transients in slow-fast systems generated by an explicit separation of timescales via a small parameter $0<\epsilon\ll 1$. The first system we consider is the FitzHugh-Nagumo model, a simplified version of the original Hodgkin-Huxley neuronal model. Here, we use the dimensionless version of the system by Cebrián-Lacasa et al. [12], given by:
$\dot{u}=u-u^{3}-v,\quad\dot{v}=\epsilon(u-bv+a),$ (5)
where $u$ represents membrane voltage of the neuron, $v$ slow recovery from depolarization through the opening of K+- and inactivation of Na+-channels, $a,b$ are dimensionless parameters, and $\epsilon$ the timescale separation. Using GhostID on a trajectory from Equation 5 in a relaxation oscillation regime with pronounced timescale separation shows again that our algorithm does not identify the slow dynamics from slow-fast systems as ghosts (Figure 3c). Further examples of GhostID correctly non-identifying long transients stemming from slow-fast dynamics rather than ghosts can be found in Supplementary section 1.4.3 and Supplementary Figure 8.
Having tested our algorithm on various models from different disciplines exhibiting different types of long transients, we conclude that GhostID correctly and specifically identifies ghost dynamics.
Figure 3: Numerical validation of GhostID. a, ghost identified by GhostID in ecological model from [8]. Parameter values: $a=2.0$, $\gamma=0.7$, $m=0.15$, $g=0.4$, $N_{t}=0.53$, $N_{0}=0.5$, $r=1.8$, $c=0.25$. b, long transients due to saddle crawl-bys in model from [27] are not identified by GhostID. Parameter values: $\gamma=2.5$, $h=1$, $v=0.5$, $m=0.4$, $\alpha=0.8$, $K=15$, $\epsilon=1$. c, long transients due to slow-fast dynamics in FitzHugh-Nagumo model are not identified by GhostID. Parameter values: $a=0$, $b=0.5$, $\epsilon=0.01$. a-c, Left: phase space. Grey dots indicate saddles, white dots indicate unstable focus. Middle: $pQ(t)$ plots along with $Q$-minima and peak prominences. Right: eigenvalues along trajectory segments and evaluation criteria.
2.4 Applications
Having defined new types of ghosts and algorithms to identify and characterize them, we seek to find examples of these phase-space objects. For this purpose, we next turn to paradigmatic models from several disciplines in which long transients and ghost dynamics have received still relatively limited attention.
2.4.1 Two-dimensional ghosts and ghost-bifurcations in coupled theta neurons
To begin our investigation into complex and potentially novel ghost dynamics, we decided to focus on simple models of coupled neurons whose dynamics in isolation exhibit a SNIC bifurcation as key ingredient for the emergence of ghosts. While long transients and metastability in computational neuroscience have received some attention over time (cf. [25, 54] for reviews), the role of ghosts in neuronal dynamics remains relatively underexplored. A good example for the importance of ghosts was recently provided by Choi et al. which proposed fly olfactory neurons to be organized close to a SNIC bifurcation, showing that the resulting ghost dynamics provide optimal information processing for sensory stimuli [13]. We thus consider the theta neuron [19] as a simple phase oscillator model that captures the essential dynamic features described in [13] to explore how pulse coupling of two identical theta neurons influences the ghost dynamics. The equations for this systems are given by [5]:
$\dot{\theta}_{1}$ $=1-\cos(\theta_{1})+\left(1+\cos(\theta_{1})\right)\left(\eta_{1}+\kappa\left(1-\cos(\theta_{2})\right)\right)$ (6)
$\dot{\theta}_{2}$ $=1-\cos(\theta_{2})+\left(1+\cos(\theta_{2})\right)\left(\eta_{2}+\kappa\left(1-\cos(\theta_{1})\right)\right),$
where $\theta_{i}$ is the phase, $\eta_{i}$ the neuron’s excitability threshold, and $\kappa$ the coupling strength determining the synaptic input current from neighboring neurons. In isolation, each theta neuron can either be excitable ($\eta<0$) or generate periodic spikes through a SNIC-bifurcation when $\eta>0$. For the coupled system we assume a coupling strength of $\kappa=0.1$. We first consider the case of two identical neurons, i.e., $\eta_{1}=\eta_{2}=\eta$. For $\eta<0$ we find the same phase space topology with four fixed points as for subsection 1.7 (Figure 4a, left). Increasing $\eta$ leads to a saddle-node bifurcation at $\eta=0$, where all four fixed points coalesce at the origin (Figure 4b). Examining this equilibrium shows that it is consistent with definition 2.1 and can be classified as a type $2,0$ saddle-node (Supplementary section 1.6). Using GhostID on a trajectory that passes through the slow region of the phase space for $\eta>0$ identifies a ghost of dimension 2 in the center of the phase space in which the flow is first funneled, then spread out again (Figure 4a, right). Note that since the phase space of Equation 6 is the torus, no stable periodic orbit exists in this topology. Using PyGhostID’s track_\_ghost_\_branch function, we can track the identified ghost as we vary the excitability threshold $\eta$ and include the ghost in the bifurcation diagram (Figure 4b). Moreover, we can extract information about each ghost along the branch and, as an example, have visually represented the trapping time at the ghosts by color, allowing us to easily identify in which parameter regime to expect significant long transients.
Figure 4: Higher-dimensional ghosts in coupled theta neuron models. a, left: phase space organization of identical theta neurons before bifurcation point. Black dot indicates sink, white dot indicates repeller, grey dots indicate saddles. Right: trajectory and two-dimensional ghost identified by GhostID after bifurcation. a,d White lines indicate trajectories, grey arrows flow. Note that for better visualization, $\theta_{1}$ and $\theta_{2}$ have been phase-shifted by $\pi$ in these plots, such that the origin is shown at $(\pi,\pi)$. b, tracking of the ghost identified in a in a bifurcation diagram using the track_\_ghost_\_branch function. Each ghost along the branch is color-coded by its trapping duration. c, two-parameter bifurcation diagram for non-identical theta neruons with distinct excitability thresholds. Boundary between areas with one- and two-dimensional ghosts indicate ghost SN-bifurcation (gSN). Red line indicates SNIC bifurcation. d, phase portraits for parameters yielding one-dimensional ghosts.
Next, we turn to the case of non-identical neurons, i.e., allowing for $\eta_{1}\neq\eta_{2}$. While track_\_ghost_\_branch currently does not feature tracking of ghosts versus two parameters, we can apply the ghostID_\_phaseSpaceSample method to characterize the ghost dynamics across the $\eta_{1}$-$\eta_{2}$-plane from a Latin hypercube sample within a specified region in phase space. Using this approach, we find two transitions from a parameter region in which only fixed points exist (i.e. without any ghosts) towards regions in which only one-dimensional ghosts exist after crossing a SNIC bifurcation (Figure 4c). Both SNIC branches merge at $\eta_{1}=\eta_{2}=0$, corresponding to two saddle-nodes of type $1,0$ coalescing at $\theta_{1}=\theta_{2}=0$ to form a type $2,0$ saddle-node. After passing one of the SNIC bifurcations, the system exhibits a stable and an unstable limit cycle due to one of the neurons still being in an excitable, not spiking regime (Figure 4d). Interestingly, moving from the 1-dimensional ghost region into the region defined by both $\eta_{1},\eta_{2}>0$ reveals another transition into a region where two-dimensional ghosts exist. This transition seems to be best described as a ghost SN-bifurcation, in which the attracting one-dimensional ghost merges with the non-attracting one-dimensional ghost shown in (Figure 4d). Indeed, we observe the same bifurcation between the ghost of a type $1,0$ saddle-node and the ghost of a type $1,1$ saddle-node for system 1.7 (Supplementary Figure 9). Together these results highlight the capability of GhostID to detect ghosts, track them across parameter space and identify transitions in their properties.
[…]
2.4.2 Timescale induced transitions in ghosts of coupled climate tipping elements
As another example, we consider coupled climate systems. While tipping of climate systems due to passing through saddle-node bifurcations is a key topic in climate research, the number of studies explicitly considering the role of ghosts is surprisingly sparse. Only recently, for example, has the potential role of ghost dynamics after tipping of the Atlantic meridian overturning circulation been taken into account [44, 9]. For studying ghost dynamics in tipping elements, we decided here to use a model of cascading climate tipping elements by Wunderling et al. [73] which is conceptually simple, yet easy to parametrize with the relevant real-world parameters and has informed a series of policy-relevant studies on climate change [74, 46]. The system is given by:
$\dot{x}_{i}=\frac{1}{\tau_{i}}\left(x_{i}-x_{i}^{3}+\sqrt{\frac{4}{27}}\frac{\Delta GMT}{T_{\text{limit},i}}+d\sum_{\forall j:j\neq i}\frac{s_{ij}}{10}(x_{j}+1)\right),$ (7)
where $x_{i}$ denotes the current value of the $i$th tipping element, $\tau_{i}$ is its tipping timescale, $\Delta GMT$ is the increase in the global mean surface temperature above pre-industrial levels, $T_{\text{limit},i}$ is the critical temperature above which the element tips (passes a SN-bifurcation), $d$ is an overall interaction strength parameter and $s_{i,j}$ represents the link strength between two interacting tipping elements. Since Equation 7 explicitly features multiple timescales, it is also an interesting system to explore how slow-fast dynamics and ghost dynamics interact. While a detailed investigation of the ghost structures in this system is the subject of another study [Nandan et al., in preparation], we focus here on two generic tipping elements which mutually reinforce each other. We assume the two tipping elements have similar but not identical critical temperatures and that their tipping timescales are different by a factor of 10. Generating several trajectories that sample the phase space using PyGhostID’s ghostID_\_phaseSpaceSample function, we identify three ghosts of dimension 1 for $\Delta GMT$ being set just above the highest value of $T_{\text{limit},i}$ (Figure 5a). Following a trajectory starting at $x_{1}(0)=-1.5,x_{2}(0)=-1.2$, we see the trajectory visiting both ghosts G3 and G1 before arriving at a stable fixed point (Figure 5a, white line), suggesting the system exhibits a ghost channel [38] between G3 and G1. To confirm this, we use the ghost_\_connections function on the ensemble of trajectories to generate an adjacency matrix of ghost-ghost sequences. In addition to plotting the corresponding graph, we can encode information about each ghost captured by GhostID visually via the node color, e.g., to represent the trapping duration at each ghost (Figure 5b). To confirm that ghostID_\_phaseSpaceSample identified all ghosts, we further followed each ghost while varying $\Delta GMT$ values using track_\_ghost_\_branch, showing that each ghost branch terminates at one of three SN-bifurcation points in which a stable fixed point and a saddle merge (Figure 5c).
Next, we explored the influence of tipping timescales in this context. We thus characterize the system’s ghost dynamics as we systematically increase $\tau_{1}$ until $\tau_{1}>\tau_{2}$. As track_\_ghost_\_branch only tracks a single ghost, we used ghostID_\_phaseSpaceSample in combination with ghost_\_connections to identify the dimensions and the connections between the ghosts. While two of the ghosts exhibited no changes, we found a transition in the dimension of G3 from 1 to 2 at $\tau_{1}\approx 1.5\ \tau_{2}$ (Figure 5d, upper panel). Interestingly, considering the number of ghost connections as quantified by the adjacency matrix reveals another transition: as $\tau_{1}$ increases, the number of ghost connections drops to 0 at $\tau_{1}=3500$ before rising to 1 again at $\tau_{1}\geq 18500$ (Figure 5d, lower panel). Visualizing the corresponding graphs reveals that these transitions are associated with a re-wiring of the ghost connections. While for low values of $\tau_{1}$, there is a ghost channel from G3 to G1, a new ghost channel from to G3 to G2 is formed for high $\tau_{1}$ values. While these two types of transitions are clearly different from each other as indicated by different critical thresholds, they are also different from the ghost SN-bifurcation described above, in which an attracting and a non-attracting ghost merge. Zooming in into the phase space region around G3, we can see how increasing $\tau_{1}$ changes both the distribution of eigenvalues across space and the direction of the flow around G3 (Figure 5e). At low $\tau_{1}$-values, we see only eigenvalue $\lambda_{1}$ around G3 being distributed from negative to positive along the flow, while $\lambda_{2}$ stays purely negative around G3. As $\tau_{1}$ increases, the flow compressed in G3 is more spread in the outbound region of G3 where now $\lambda_{2}$, too, becomes positive, thereby allowing trajectories to sample $\lambda_{1}$ from negative to positive. Hence, the change in the dimension of G3 is as result of both the changed eigenvalue distribution across space and of the altered flow, which together determine what eigenvalues a trajectory will see as it passes close to or through G3. The transition between ghost connections, on the other hand, is a result only of the altered flow that determines if a another ghost can be reached starting from G3 or not.
2.4.3 Ghost networks in gene regulatory networks at criticality
As a last example we consider ghost dynamics in GRNs in which each gene exhibits positive auto-regulation and can further be activated and repressed by other genes, modeled in an additive and multiplicative fashion by Hill-equations [2] with a Hill-coefficient of 2, respectively:
$\dot{x_{i}}=(\frac{\alpha x_{i}^{2}}{x_{i}^{2}+K^{2}}+\sum_{\forall j:a_{ji=1}}\frac{\beta x_{j}^{2}}{x_{j}^{2}+K_{a}^{2}})\prod_{\forall j:a_{ji=-1}}\frac{K_{i}^{2}}{x_{j}^{2}+K_{i}^{2}}-x_{i},$ (8)
where $x_{i}$ refers to the concentration of gene product $i$, $\alpha$ is the auto-activation strength, $\beta$ the maximum transcription rate of activators, $a_{ij}$ are the elements of the adjacency matrix $A$ that describes the GRN’s interactions, and $K,K_{a},K_{i}$ are the half-activation constants of (auto-)activation and repression, respectively. For $K=0.5$ and in the absence of any activators and inhibitors, each gene exhibits a SN-bifurcation at $\alpha=1$ (Figure 6a). We assume here $\alpha=0.998$, i.e., each gene is organized close to the SN-bifurcation and may thus exhibit a simple ghost, not unlike the situation described in [20]. Directed Erdős–Rényi networks for which each activatory link was replaced with an inhibitory link with a probability of $0.5$ were used to generate different GRN topologies.
Using this simple network model, we observed larger ghost structures including ghost channels, ghost cycles and various combinations thereof, with many ghosts having a dimension larger than one. As an example we show a network with $N=12$ nodes (Figure 6b). Using a large phase space sample of this network generated by ghostID_\_phaseSpaceSample, we identified the ghost structures embedded in the GRN using the ghost_\_connections function and colored each node in the corresponding graph according to the ghost dimensions. As shown in Figure 6c, we find a strikingly complex ghost network embedded in the phase space of the GRN, featuring many interesting motifs, including cyclical structures, several ghosts converging into one, and ghosts ejecting their flow to two or more other ghosts (e.g. G6). While a systematic characterization of such ghost networks in phase space and their relation to the network structure of the dynamical system is beyond the scope of the current study, a few interesting observations can be drawn from this example already. First, the ghost network has a different topology and is notably larger than the GRN, featuring both more nodes and edges. The topology of the ghost network further suggests multi-stability between a three-ghost cycle, at which the majority of nodes eventually converge by following their links, and two ghost channels from G5 and G9 to G10 after which trajectories converge at an attractor. Timecourses of trajectories starting at three random initial conditions confirm this, showing convergence to the three ghost cycle with its periodic switching between distinct long transients of gene expression after different initial transients for two initial conditions and a two-ghost channel that ends at a stable steady state for the third initial condition (Figure 6d). Another interesting observation is that the dimensions of the ghosts in the ghost network are relatively small compared to the dimension of the system ($n=12$), with the majority of ghosts having a dimension of 2-4. The dimension of the ghosts furthermore do not necessarily correspond to the number of connections a ghost forms. For example, while G9 has a ghost dimension of four, it only connects to two other ghosts. Sampling initial conditions around this ghost and projecting the high-dimensional trajectories into a three-dimensional principle component space shows that trajectories are being spread into four different directions, consistent with the ghost dimension measured by GhostID, before two pairs of the four trajectory bundles are each being funneled towards G4 and G10, respectively (Figure 6e).
Figure 6: Complex ghost structures in gene regulatory networks. a, bifurcation diagram for single gene dynamics described by equation 8. b, topology of a GRN generated by the Erdős–Rényi random model. c, toplogy of a complex ghost network embedded in the dynamics of the GRN shown in b. Network was identified from a large phase space sample (300 trajectories) using the ghostID_\_phaseSpaceSample and ghost_\_connections functions from PyGhostID. d, timecourses of the GRN shown in b for three different random initial conditions. e, trajectories shown in principal component space for multiple initial conditions starting close to G9.
3 Discussion
We have presented here a generalized definition of ghosts in dynamical systems that accounts for the diverse topological structures of the saddle-node fixed points they are the remnant of and developed a series of algorithms to identify and characterize ghosts and their composite structures such as ghost channels and ghost cycles [38]. Naturally, there are several open questions and limitations that come with our approach. First, our current framework was developed to capture ghosts of fixed points. While further work is necessary to see whether it applies to ghosts of continuous attractors [23, 56], the current algorithms cannot capture ghosts of limit cycles as the slow dynamics only apply along the radial direction of their periodic orbit and it is unclear if any meaningful eigenvalue characteristics can be extracted from ghosts of limit cycles. The same argument applies to ghosts from chaotic attractors [44, 9]. At least the first problem has been solved by the elegant study from Zheng et al. [77], who used variational methods to create a ‘best-fit’ topological structure in the vector field. An alternative approach may be to replace the scalar function $Q$ by recurrences [15] for finding non-invariant structures in phase space that a trajectory spends a long time in. Secondly, the framework was designed for autonomous deterministic systems. While we suspect that slow parameter drifts will not impede detection of ghosts by our algorithm, faster or more complex parameter changes in non-autonomous systems will likely interfere with the algorithm. Similar considerations apply to stochastic systems. Noise has a significant influence on a trajectories’ dwell-time at a ghost and can push it back and forth along the ghost region [57, 38] which may result in non-monotonic eigenvalue samples that impair ghost identification. Lastly, we assumed throughout that $f$ is analytic. While this assumption holds for a vast number of models from applications, additional assumptions are likely required for our framework to apply to smooth systems (cf. Supplementary section 1.7), although these assumptions are unlikely to affect the here introduced algorithms in most cases.
Despite these limitations, we showed that our definitions and algorithms offer a powerful new approach to study ghost dynamics in arbitrary dynamical systems. Using several relatively simple example systems, this approach allowed us to discover several interesting phenomena related to ghosts. While their comprehensive characterization is beyond the scope of the current study, we will briefly discuss their implications.
In our first example, two coupled and identical theta neurons [5], we showed the existence of a two-dimensional ghost following the merging of a sink, a repeller and two saddles (a phase portrait sometimes called basic tartan [6]) at a type $2,0$ saddle-node of a degenerate SNIC-bifurcation. Since the theta neuron model is the normal form of a SNIC bifurcation, there is strong reason to believe that this result will apply to many coupled systems whose units individually undergo SNIC bifurcations. Characterizing the ghost dynamics in the case of non-identical neurons with different excitability thresholds, we found that one- and two-dimensional ghosts are detectable across a large area of the parameter plane, although pronounced long transients are only found close to bifurcation. This two-parameter analysis further revealed the existence of what we termed a ghost SN-bifurcation, a collision of two one-dimensional ghosts - one of which is attracting, one non-attracting in the sense of definition 2.2 - leading to the formation of a two-dimensional ghost. To our knowledge, this is the first description of a bifurcation that occurs exclusively between non-invariant objects and produces another non-invariant object. From a biological perspective, the example raises interesting new questions about the role of ghost dynamics in the context of information processing and computations in neuronal networks. For example, could higher dimensional ghosts provide a reliable mechanism for decoding multi-modal time-varying inputs as opposed to single time-varying inputs considered previously [13]? Assuming larger ghost networks as found in our GRN example can also emerge from neuronal dynamics, another intriguing perspective is to consider ghosts and their composite structure as a potential transient-based mechanism for structuring phase space trajectories along neuronal manifolds [50]. As we find the properties of ghosts to be tunable by control parameters, it is tempting to speculate that neuronal networks might be able to adapt the properties of individual ghosts and larger ghost structures to their computational needs. Studying how ghost dynamics structure and support biological computations is thus an important area for further research.
In the second example, we examined coupled climate tipping elements using the model from Wunderling et al. [73]. We focused here on two the case of two generic, mutually reinforcing tipping elements with similar temperature thresholds but different tipping timescales in a scenario when the global mean temperature just passed these thresholds. Our analysis revealed the existence of three ghost states in this model, two of which form a ghost channel, detectable for the whole range of global warming considered here (i.e., up to $\Delta GMT=3$°C). We further found a striking influence of the tipping timescales on the system’s ghost dynamics, which included the change of one of the ghosts dimension by altering the local flow and eigenvalue distribution as well as a restructuring of the ghost connections. Since neither of these transition involved a collision between ghosts, the observed transitions are distinct from the ghost SN-bifurcation observed in the coupled theta neuron model. However, further work is required to better understand the nature of the timescale-dependent transitions. Since long transients due to ghosts can keep a system in a desirable area of phase space even when the equilibrium is lost, a more detailed investigation of the ghost dynamics in a higher dimensional version of this model for realistic parameters, real-world network topologies and policy scenarios is currently underway [Nandan et al., in preparation].
In the final example, we applied our algorithms to GRNs under the assumption that each gene is poised at criticality, i.e., close to a SN-bifurcation. Our analysis showed that embedded in the dynamics of such GRNs can be a complex ghost network in phase space with many higher-dimensional ghosts and a topology distinct from the GRN. These ghost networks can further be multi-stable in the sense of coexisting ghost structures, such that some initial conditions lead to visitation, e.g., of a ghost channel, whereas others end up on a ghost cycle. While the assumption of each gene being poised at criticality is perhaps too strong to be representative of real GRNs, there is a growing body of empirical evidence for ghost dynamics in GRNs [53, 45] and intracellular signaling networks [62]. While the role of ghost dynamics in cell biology is barely explored, its unique properties have been used to explain processing of time-varying stimuli[63, 37] and put forward as robust mechanisms for cell population control[61], regulation of developmental transitions [53], and cell-type specification and differentiation [45, 20]. Moreover, our example demonstrates that complex ghost dynamics such as described in [20] can easily emerge in random networks and do not require artificial network structures. Exploring complex and higher-dimensional ghost dynamics in GRNs and other types of intracellular regulatory networks thus has the potential to provide new perspectives on various phenomena in cell biology.
Based on these results, we believe the framework and algorithms presented here will open many interesting avenues for further research ranging from physical sciences to living systems.
4 Methods
All models in this study were implemented in Python and simulations were performed using the RK45 integration scheme as implemented in SciPy’s solve_\_ivp function [71]. Numerical bifurcation analysis of the coupled theta neurons and tipping element models was performed using XPPAUT. A detailed description of the GhostID algorithm for identifying ghosts from trajectories is given in subsection 2.2 and Supplementary Figure 2. Algorithms for phase space sampling, continuing identified ghosts across parameters and identifying composite ghost structures such as ghost channels and ghost cycles are described in Supplementary section 1.2.
Author contributions
D.K.: conceptualization, formal analysis, numerical simulations, writing—original draft. A.P.N.: formal analysis, writing—review & editing.
Competing Interests
The authors declare no competing interests.
Use of generative AI statement
Generative AI has been used for specific coding tasks and to check the validity of the introduced definitions and theorems. A detailed overview of its use can be found in Supplementary section 3.
Acknowledgments
We would like to thank Gayathri Ramesan for helpful discussions during the early stages of this work and Aneta Koseska and Peter Ashwin for their feedback on this manuscript.
Data availability
All codes for reproducing the data and figures from this study are available at https://github.com/KochLabCode/PyGhostID. PyGhostID is available at https://pypi.org/project/PyGhostID/.